OpenStudy (anonymous):

3/m-4 + 1/3(m-4) = 6/m The only other info it gives says: You can solve a rational equation by multiplying each side by the LCD. Check possible solutions to make sure each satisfies the original equation.

OpenStudy (johnweldon1993):

\[\large \frac{3}{m - 4} + \frac{1}{3}(m - 4) = \frac{6}{m}\] Like that right?

OpenStudy (johnweldon1993):

Or is it \[\large \frac{3}{m - 4} + \frac{1}{3(m - 4)} = \frac{6}{m}\]

OpenStudy (johnweldon1993):

Ehh for simplicity I believe it would be the second one I posted... So what is the LCD for that one? Well if you see...the only difference between the 2 denominators....is the second one is the first one..just multiplied by 3....so lets multiply the whole first fraction by 3/3 to even that out... \[\large \frac{3}{3} \times \frac{3}{(m - 4)} + \frac{1}{3(m - 4)} = \frac{6}{m}\] \[\large \frac{9}{\color \red{3(m - 4)}} + \frac{1}{\color \red{3(m - 4)}} = \frac{6}{m}\] Now we have a common denominator ...so lets combine those 2 fractions over that common denominator \[\large \frac{9 + 1}{3(m - 4)} = \frac{6}{m}\] \[\large \frac{10}{3(m - 4)} = \frac{6}{m}\] Lets cross multiply here \[\large 10m = 6 \times 3(m - 4)\] \[\large 10m = 18(m - 4)\] Can you go from there?

OpenStudy (anonymous):

@DollyAcquah he is offline, could you help me finish this?

OpenStudy (anonymous):

sorry to far ahead of me im in 8th grade

OpenStudy (anonymous):

@natedog12 @secret66 @Luigi0210 @ShaTay @YanaSidlinskiy

OpenStudy (anonymous):

@wyattp17

OpenStudy (anonymous):

do you have any answer choices?

OpenStudy (yanasidlinskiy):

Does @johnweldon1993 not help you?? I promise...He will get you through it...

OpenStudy (anonymous):

not me

OpenStudy (anonymous):

i have the answer i just need to show my work

OpenStudy (anonymous):

he is not online anymore! but he was helping alot!

OpenStudy (anonymous):

the answer is 9 but i need to show my work, and i dont know how to do this question

OpenStudy (yanasidlinskiy):

Trust me...Either message or email him..

OpenStudy (anonymous):

well if he is offline he isnt going to get it until he comes back online.. but i wanted to get this done soon. I just needed to know how to show the work for it. @geniousmind13 the answer is 9, do you know the work to show for that?

OpenStudy (anonymous):

yes just a minute and i will have it posted

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

here it is hope it helps

OpenStudy (johnweldon1993):

@Reilly208 still online??

OpenStudy (johnweldon1993):

Well I'm wondering if you understood everything I have done so far...

OpenStudy (anonymous):

hi john gues what

OpenStudy (johnweldon1993):

If so... we stopped after cross multiplying and now we have \[\large 10m = 18(m - 4)\] Distribute that 18 into the parenthesis \[\large 10m = 18m - 72\] Combine like terms...subtract 18m from both sides of this equation \[\large -8m = -72\] And finally just divide both sides by -8 \[\large m = 9\]

OpenStudy (johnweldon1993):

what? @natedog12

OpenStudy (anonymous):

my dad names is john